In two years time, Ella's age will be three times Lucy's age. Lucy is now t years old. How old is Ella now?
Answer by stanbon(75887)

In two years time, Ella's age will be three times Lucy's age. Lucy is now t years old. How old is Ella now?
-------------
Equations:
e + 2 = 3(L+2)
L = t
-----
e + 2 = 3L + 6
Ella = 3t +4
======================
Cheers,
Stan H.